### v2

# Spectral Broadening of Landau Levels by a Penetrable Circular Wall

Abstract
We study the two–dimensional magnetic Schrödinger operator with a penetrable circular wall modeled by a $\delta$–interaction. Using the boundary triple approach we classify all self–adjoint extensions and obtain Krein’s resolvent formula, showing that the essential spectrum coincides with the Landau levels. The wall breaks their infinite degeneracy and produces a spectral broadening: each Landau level becomes an accumulation point of discrete eigenvalues from one side. In the circular case, rotational symmetry reduces the eigenvalue problem to scalar equations with explicit Weyl coefficients. We prove strict monotonicity, ensuring that each angular momentum channel contributes at most one eigenvalue per gap, and derive asymptotics showing that the boundary coefficients decay faster than any exponential, explaining the strong localization of the broadened spectrum. Numerical simulations are consistent with these results.
1 Introduction
Landau levels are a characteristic feature of quantum mechanics in a magnetic field. They represent an infinite degeneracy of energy levels, reflecting the high symmetry of cyclotron motion. Understanding how this delicate structure is modified under perturbations has been a central theme in the spectral theory of magnetic Schrödinger operators.
In the planar case with a constant magnetic field $B>0$, the free Hamiltonian
$H_{B,0}=(-i\nabla-A)^{2}$ (1.1)
has spectrum consisting only of the infinitely degenerate eigenvalues
$\Lambda_{n}=B(2n+1),\qquad n\in\mathbb{N}_{0},$ (1.2)
known as the Landau levels; see, for instance, [15]. It is well known that under various perturbations, discrete eigenvalues may appear close to these levels (cf. [18, 16, 7]).
Among the most natural perturbations are singular interactions supported on lower-dimensional sets. $\delta$–type interactions provide models of thin barriers or penetrable–walls and admit a rigorous operator-theoretic treatment[1, 2, 11]. Even without a magnetic field, such interactions can produce curvature-induced bound states [9, 10]. In the magnetic case, $\delta$–interactions supported on general curves were shown to break the degeneracy of Landau levels and to produce discrete eigenvalues accumulating near them [3]. These results, however, are mainly qualitative: they establish the existence of clusters but do not give a detailed description of their internal structure.
The aim of this paper is to present a detailed and explicit analysis of the two-dimensional Landau Hamiltonian with a penetrable circular wall, that is, a $\delta$–interaction supported on the circle $S_{a}=\{x\in\mathbb{R}^{2}:|x|=a\}$. Formally, the model can be written as
$H_{B,\alpha}=H_{B,0}+\alpha(x)\delta(|x|-a),$ (1.3)
where $\delta(|x|-a)$ denotes the Dirac interaction supported on $S_{a}$. A rigorous definition of $H_{B,\alpha}$ will be given later in the framework of boundary triples.
By Weyl’s theorem, the Landau levels $\Lambda_{n}$ remain in the essential spectrum for $\alpha\neq 0$. However, they are no longer eigenvalues of infinite multiplicity; instead, each $\Lambda_{n}$ becomes an accumulation point of infinitely many finite-multiplicity eigenvalues created by the wall.
In the circular geometry, rotational symmetry reduces the eigenvalue problem to scalar equations involving the diagonal coefficients $\mu_{m,B}(z)$ of the Weyl function. We obtain explicit formulas for these coefficients (Lemma 11) and prove their strict monotonicity (Proposition 12), showing that each angular momentum channel contributes at most one eigenvalue in each spectral gap. An asymptotic analysis of the boundary coefficients $c_{n,m}$ (Proposition 16) shows that they decay faster than any exponential in $|m|$, which explains the strong localization of the broadened spectrum. Numerical simulations are consistent with these results. We restrict ourselves here to the two-dimensional case, where genuine spectral broadening of Landau levels occurs.
In addition to these analytic results, the circular wall is not only the simplest nontrivial geometry that allows a fully explicit treatment, but also provides new physical insight: the infinite degeneracy of Landau levels is decomposed into resonant angular momentum modes, whose interaction with the boundary governs the fine spectral structure. This mechanism, not visible in the abstract theory, shows how geometry influences localization in magnetic quantum systems.
Novelty compared with previous work. The general theory developed in [3] shows the existence of eigenvalue clusters near Landau levels for $\delta$–interactions supported on curves of arbitrary shape, but the results are essentially qualitative. By focusing on the circular geometry, we obtain explicit and quantitative information on the fine structure of the broadened spectrum:
  • (i)
    An explicit diagonalization of the Weyl function, reducing the eigenvalue condition to the scalar equations $\alpha-\mu_{m,B}(E)=0$ (see Proposition 6).
  • (ii)
    A strict monotonicity result (Proposition 12) showing that each angular momentum channel contributes at most one eigenvalue per spectral gap.
  • (iii)
    Precise asymptotics for the boundary coefficients $c_{n,m}$ (Proposition 16), establishing faster than exponential decay in $|m|$, and the corresponding asymptotic formula for the eigenvalue shifts (Theorem 17).
  • (iv)
    A semiclassical interpretation (Section 3.6) that links the wall radius with resonant cyclotron modes and explains the mode selectivity in geometric terms, a viewpoint not available in the abstract theory.
Taken together, these results show that the essential spectrum remains the Landau levels, while for each $n$ the operator $H_{B,\alpha}$ exhibits an infinite sequence of discrete eigenvalues that accumulate at $\Lambda_{n}$ from one side only, with the sign of $\alpha$ determining whether the accumulation occurs from above or from below. These analytic findings are consistent with numerical simulations, implemented in Python, which provide an intuitive visualization of the spectral broadening phenomena discussed in this work.
2 Self–adjoint realizations in two dimensions
In this section we construct the minimal magnetic operator with a circular interface and establish the associated Green’s identity. Using the boundary triple approach we classify all self–adjoint extensions and introduce the penetrable–wall Hamiltonians, which will serve as the basis for the spectral analysis in the subsequent sections (cf. [12, 6, 5, 17, 2]).
We note that in the present two–dimensional magnetic setting the boundary maps naturally yield a quasi boundary triple in the sense of [4, 17].
2.1 Comparison with the three–dimensional case
We recall the classical three–dimensional model studied by Ikebe and Shimada [13], where the Laplacian in $\mathbb{R}^{3}$ is coupled to a penetrable spherical wall at radius $a>0$. The corresponding Hamiltonian is defined by
$H_{\alpha}^{3D}u$ $=-\Delta u,$
$\mathcal{D}(H_{\alpha}^{3D})$ $=\Bigl{\{}\,u=(u_{-},u_{+})\in H^{2}(\Omega_{-})\oplus H^{2}(\Omega_{+}):$
$\qquad u_{+}|_{S_{a}}=u_{-}|_{S_{a}},\qquad[\partial_{r}u]=\alpha\,u|_{S_{a}}\,\Bigr{\}}.$
where $S_{a}=\{x\in\mathbb{R}^{3}:|x|=a\}$ denotes the sphere (we use the same notation as for the circle in two dimensions), $\Omega_{-}=\{x\in\mathbb{R}^{3}:|x|<a\,\}$ is the interior ball, and $\Omega_{+}=\{x\in\mathbb{R}^{3}:|x|>a\,\}$ is the exterior domain. Here $\partial_{r}$ denotes the radial derivative, and $\alpha\in\mathbb{R}$ is the coupling strength. The jump of the radial derivative across the sphere is defined, for $u=(u_{-},u_{+})$ with $u_{\pm}\in H^{2}(\Omega_{\pm})$, by
$\partial_{r}u]\;:=\;\partial_{r}u\big{|}_{r=a^{+}}\;-\;\partial_{r}u\big{|}_{r=a^{-}}.$
They proved that the essential spectrum of $H_{\alpha}^{3D}$ is $0,\infty)$, and that discrete eigenvalues may appear below 0 depending on $a$ and $\alpha$. In particular, in the attractive case $\alpha<0$ there may exist finitely many negative eigenvalues, while in the repulsive case $\alpha>0$ no discrete spectrum occurs (and, in fact, no embedded eigenvalues appear in $0,\infty)$). The occurrence of negative eigenvalues in the attractive case is linked to the radial part of the problem, governed by spherical Bessel functions, at the interface $r=a$.
In contrast, the two–dimensional model studied in the present paper exhibits qualitatively different behavior. The free Hamiltonian with magnetic field $B>0$ has spectrum given by the Landau levels (1.2), each of which is an eigenvalue of infinite multiplicity. Introducing a penetrable circular wall at $S_{a}$ breaks this degeneracy, but instead of producing finitely many isolated eigenvalues, one obtains infinite sequences of eigenvalues accumulating at each $\Lambda_{n}$. In other words, each Landau level becomes an accumulation point of a one–sided spectral broadening, a phenomenon that has no counterpart in the three–dimensional case.
This fundamental difference stems from the interplay between magnetic quantization and two–dimensional geometry: in 2D the Landau levels act as highly degenerate accumulation points, and the wall interaction splits this degeneracy into angular momentum channels, each contributing at most one eigenvalue per gap. The accumulation of these eigenvalues at $\Lambda_{n}$ constitutes the spectral broadening analyzed in detail in Sections 35.
2.2 Preliminaries and the minimal operator
Let $B>0$ and fix a smooth vector potential $A:\mathbb{R}^{2}\to\mathbb{R}^{2}$ with $\operatorname{rot}A=B$. For definiteness we may choose the symmetric gauge $A(x)=\tfrac{B}{2}(-x_{2},x_{1})$, although the following constructions are in fact gauge invariant. The free magnetic Laplacian is then defined by (1.1), acting in $L^{2}(\mathbb{R}^{2})$ with initial domain $C_{0}^{\infty}(\mathbb{R}^{2})$. Let $a>0$ and denote by $S_{a}=\{x\in\mathbb{R}^{2}:\ |x|=a\}$ the boundary circle. We write $\Omega_{-}=\{\,|x|<a\,\}$ and $\Omega_{+}=\{\,|x|>a\,\}$, with unit normal $\nu$ pointing from $\Omega_{-}$ to $\Omega_{+}$.
In order to incorporate the wall, we consider the minimal operator with an interface at $S_{a}$,
$Tu:=(-i\nabla-A)^{2}u,\qquad\mathcal{D}(T):=\overline{C_{0}^{\infty}(\mathbb{R}^{2}\!\setminus\!S_{a})}^{\|\cdot\|_{L^{2}}+\|H_{B,0}\cdot\|_{L^{2}}},$ (2.1)
which is closed and symmetric in $L^{2}(\mathbb{R}^{2})$ (see, e.g., [5] for the elliptic operator framework). For $u\in H^{2}(\Omega_{+})\oplus H^{2}(\Omega_{-})$ we set the magnetic normal derivatives
$\partial_{\nu}^{A}u_{\pm}:=\nu\cdot(\nabla-iA)\,u_{\pm}\big{|}_{S_{a}}.$
We also introduce the average trace and the jump of magnetic normal derivatives by
$\gamma_{\mathrm{av}}u:=\tfrac{1}{2}\big{(}\gamma^{+}u_{+}+\gamma^{-}u_{-}\big{)},\qquad[\partial_{\nu}^{A}u]:=\partial_{\nu}^{A}u_{+}-\partial_{\nu}^{A}u_{-},$
where $\gamma^{\pm}$ are the standard Sobolev traces. With these notations one verifies the Green’s identity, valid for all $u,v\in H^{2}(\Omega_{+})\oplus H^{2}(\Omega_{-})$,
$\langle(-i\nabla-A)^{2}u,v\rangle_{L^{2}(\Omega_{+})}-\langle u,(-i\nabla-A)^{2}v\rangle_{L^{2}(\Omega_{+})}$ (2.2)
$\quad+\langle(-i\nabla-A)^{2}u,v\rangle_{L^{2}(\Omega_{-})}-\langle u,(-i\nabla-A)^{2}v\rangle_{L^{2}(\Omega_{-})}$
$=\langle[\partial_{\nu}^{A}u],\,\gamma_{\mathrm{av}}v\rangle_{H^{-1/2},H^{1/2}}-\langle\gamma_{\mathrm{av}}u,\,[\partial_{\nu}^{A}v]\rangle_{H^{1/2},H^{-1/2}}.$
see, for instance, standard references on elliptic regularity and trace mappings [21]. The subsequent boundary triple formulation and the Krein’s resolvent formula we employ below follow the general scheme in [12, 6, 5, 17]; for singular perturbations modeled by $\delta$–type terms we also refer to [1, 2].
2.3 Boundary operators and self–adjoint extensions
Recall the minimal operator $T$ defined in (2.1). The boundary operators are now defined on $\mathcal{D}(T^{\ast})=H^{2}(\Omega_{+})\oplus H^{2}(\Omega_{-})$ by
$\Gamma_{0}u:=[\partial_{\nu}^{A}u],\qquad\Gamma_{1}u:=\gamma_{\mathrm{av}}u.$
Then the triple $\big{(}H^{1/2}(S_{a}),\Gamma_{0},\Gamma_{1}\big{)}$ forms a quasi boundary triple for $T^{\ast}$ (in the sense of the standard theory, see [12, 6, 4]). In particular, the abstract Green’s identity (2.2) holds for all $u,v\in\mathcal{D}(T^{\ast})$. This framework will serve as the basis for the Krein resolvent formula in the standard form
$(H_{B,\Theta}-z)^{-1}=(H_{B,0}-z)^{-1}+\gamma_{B}(z)\,(\Theta-M_{B}(z))^{-1}\,\gamma_{B}(\bar{z})^{\ast},$ (2.3)
which we adopt consistently throughout this paper (cf. [6, 17, 4]).
Definition 1 (Weyl function).
For $z\in\mathbb{C}\setminus\mathbb{R}$, we define the Weyl function by
$M_{B}(z):=\tau(H_{B,0}-z)^{-1}\tau^{\ast},$
acting from $H^{1/2}(S_{a})$ to $H^{-1/2}(S_{a})$, where $\tau$ denotes the trace to $S_{a}$. This coincides with the abstract Weyl function associated with the boundary triple $\big{(}H^{1/2}(S_{a}),\Gamma_{0},\Gamma_{1}\big{)}$; see Proposition 10.
The mapping properties of $\tau$ and $\tau^{\ast}$, together with the canonical injection $H^{1/2}(S_{a})\hookrightarrow H^{-1/2}(S_{a})$ via the $L^{2}(S_{a})$–pivot, show that the Weyl function $M_{B}(z)$ acts as a bounded operator $H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$. This is summarized by the following diagram. Here $\tau:H^{2}(\mathbb{R}^{2})\to H^{3/2}(S_{a})$ is the trace map, and its adjoint $\tau^{\ast}:H^{-3/2}(S_{a})\to H^{-2}(\mathbb{R}^{2})$. By duality and the Sobolev embeddings
$H^{1/2}(S_{a})\hookrightarrow H^{-1/2}(S_{a})\hookrightarrow H^{-3/2}(S_{a}),\qquad L^{2}(\mathbb{R}^{2})\hookrightarrow H^{-2}(\mathbb{R}^{2}),$
$\tau^{\ast}$ extends canonically to a bounded operator $H^{-1/2}(S_{a})\to H^{-2}(\mathbb{R}^{2})$. Moreover, since $(H_{B,0}-z):H^{2}(\mathbb{R}^{2})\to L^{2}(\mathbb{R}^{2})$ is an isomorphism for $z\in\mathbb{C}\setminus\mathbb{R}$, its inverse extends to a bounded operator $(H_{B,0}-z)^{-1}:H^{-2}(\mathbb{R}^{2})\to L^{2}(\mathbb{R}^{2})$. Thus the composition can be understood diagrammatically as
$H^{1/2}(S_{a})$ $\hookrightarrow\ H^{-1/2}(S_{a})\ \xrightarrow{\ \tau^{\ast}\ }\ H^{-2}(\mathbb{R}^{2})$ (2.4)
$\xrightarrow{\ (H_{B,0}-z)^{-1}\ }\ L^{2}(\mathbb{R}^{2})\ \xrightarrow{\ \tau\ }\ H^{3/2}(S_{a})\ \hookrightarrow\ H^{-1/2}(S_{a}).$
Hence $M_{B}(z)$ indeed maps $H^{1/2}(S_{a})$ continuously into $H^{-1/2}(S_{a})$.
Proposition 1.
Let $\{H^{1/2}(S_{a}),\Gamma_{0},\Gamma_{1}\}$ be the (quasi) boundary triple fixed above, and let $\Theta:H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$ be bounded and self–adjoint. Assume that there exists $z_{0}\in\mathbb{C}\setminus\mathbb{R}$ such that $\Theta-M_{B}(z_{0})$ is boundedly invertible on $H^{1/2}(S_{a})$. Define
$D(H_{B,\Theta})=\{u\in D(T^{\ast}):\ \Gamma_{0}u=\Theta\,\Gamma_{1}u\},\qquad H_{B,\Theta}u=T^{\ast}u.$
Then $H_{B,\Theta}$ is self–adjoint in $L^{2}(\mathbb{R}^{2})$. Moreover, for every $z\in\mathbb{C}\setminus\mathbb{R}$ such that $\Theta-M_{B}(z)$ is invertible, one has the Krein’s resolvent formula
$(H_{B,\Theta}-z)^{-1}=(H_{B,0}-z)^{-1}+\gamma_{B}(z)\,(\Theta-M_{B}(z))^{-1}\,\gamma_{B}(\bar{z})^{\ast}.$ (2.5)
In particular, (2.3) holds for all $z\in\rho(H_{B,\Theta})\cap\rho(H_{B,0})$.
Remark 1.
We work within the quasi boundary triple framework; see, e.g., [4, 6, 17]. The sign convention in (2.3) corresponds to $(\Gamma_{0},\Gamma_{1})=([\partial_{\nu}^{A}\cdot],\,\gamma_{\mathrm{av}}\cdot)$ and avoids ambiguities among alternative formulas in the literature. The assumption that $\Theta-M_{B}(z_{0})$ is invertible for some nonreal $z_{0}$ is standard and ensures that the extension defined by $\Gamma_{0}u=\Theta\,\Gamma_{1}u$ is closed and self–adjoint in $L^{2}(\mathbb{R}^{2})$.
Proposition 2.
The deficiency indices of $T$ are infinite. In particular, $T$ is not essentially self–adjoint and admits infinitely many self–adjoint extensions.
Proof.
Let $z\in\mathbb{C}\setminus\mathbb{R}$ be fixed. By the boundary triple construction recalled in Section 2 (see also [12, 6]), for every $\varphi\in H^{-1/2}(S_{a})$ there exists a unique $\gamma_{B}(z)\varphi\in\ker(T^{\ast}-z)$ such that
$(T^{\ast}-z)\,\gamma_{B}(z)\varphi=0,\qquad\Gamma_{0}\big{(}\gamma_{B}(z)\varphi\big{)}=\varphi,$
and the Poisson operator $\gamma_{B}(z):H^{-1/2}(S_{a})\to L^{2}(\mathbb{R}^{2})$ is bounded. In particular, $\Gamma_{0}:\ker(T^{\ast}-z)\to H^{-1/2}(S_{a})$ is onto and
$\ker(T^{\ast}-z)=\mathrm{ran}\,\gamma_{B}(z).$
Since $\Gamma_{0}\circ\gamma_{B}(z)=\mathrm{Id}$ on $H^{-1/2}(S_{a})$, the operator $\gamma_{B}(z)$ is injective. Hence
$\dim\ker(T^{\ast}-z)=\dim H^{-1/2}(S_{a}).$
The space $H^{-1/2}(S_{a})$ is infinite dimensional, therefore the deficiency space $\ker(T^{\ast}-z)$ is infinite dimensional. Taking $z=i$ and z=iz=-i gives that both deficiency indices of $T$ are infinite. In particular, $T$ is not essentially self–adjoint.
Finally, by the quasi boundary triple parametrization of self–adjoint extensions (see [6, 5, 4, 17]), if $\Theta:H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$ is bounded and self–adjoint and there exists $z_{0}\in\mathbb{C}\setminus\mathbb{R}$ such that $\Theta-M_{B}(z_{0})$ is boundedly invertible, then the extension defined by $\Gamma_{0}u=\Theta\,\Gamma_{1}u$ on $S_{a}$ is self–adjoint in $L^{2}(\mathbb{R}^{2})$. Since the boundary space is infinite dimensional, there are infinitely many such $\Theta$, and hence infinitely many self–adjoint extensions.
$\Box$
Moreover, for $z\in\mathbb{C}\setminus\mathbb{R}$, the map $\Gamma_{0}:\ker(T^{\ast}-z)\to H^{-1/2}(S_{a})$ is onto, so the Poisson operator $\gamma_{B}(z)$ and the Weyl function $M_{B}(z)$ are well defined. The boundary triple formalism yields a complete parametrization of self-adjoint extensions: for any bounded self-adjoint $\Theta:H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$, the operator $H_{B,\Theta}$ defined by $\Gamma_{0}u=\Theta\,\Gamma_{1}u$ on $S_{a}$ is self-adjoint. Furthermore, $M_{B}$ is a Herglotz (Nevanlinna) function:
$\frac{M_{B}(z)-M_{B}(\bar{z})^{\ast}}{z-\bar{z}}=\gamma_{B}(\bar{z})^{\ast}\gamma_{B}(z)\ \geq 0,\qquad z\in\mathbb{C}\setminus\mathbb{R},$
that is, $M_{B}(\bar{z})=M_{B}(z)^{\ast}$ and, for $\text{{\rm Im}}z>0$,
$\text{{\rm Im}}M_{B}(z)=\frac{1}{2i}\big{(}M_{B}(z)-M_{B}(z)^{\ast}\big{)}\ \geq 0\quad\text{on }H^{1/2}(S_{a}).$
2.4 Penetrable–wall Hamiltonians and gauge covariance
We write $\mathcal{M}_{\alpha}$ for the multiplication operator by $\alpha$ to avoid confusion with the Weyl function $M_{B}(z)$, and recall from (2.6) that $\mathcal{M}_{\alpha}$ acts as a bounded operator $H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$ defined by
$\langle\mathcal{M}_{\alpha}\phi,\psi\rangle_{H^{-1/2},H^{1/2}}=\int_{S_{a}}\alpha\,\phi\,\overline{\psi}\,d\sigma\qquad(\phi,\psi\in H^{1/2}(S_{a})),$ (2.6)
where $d\sigma$ denotes the surface measure on $S_{a}$. In the following we shall simply write $\alpha$ in place of $\mathcal{M}_{\alpha}$ when no confusion can arise, since from Section 3 onwards $\alpha$ is always a constant. In this case the corresponding self–adjoint extension will be denoted $H_{B,\alpha}$. Its domain can be described explicitly as
$\mathcal{D}(H_{B,\alpha})$ $=$ $\Bigl{\{}\,u=(u^{-},u^{+})\in H^{2}(\Omega^{-})\oplus H^{2}(\Omega^{+}):$ (2.7)
$\qquad\gamma^{+}u=\gamma^{-}u=:\gamma u,\;[\partial_{\nu}^{A}u]=\alpha\gamma u\ \text{on }S_{a}\,\Bigr{\}}.$
and on this domain the operator acts as $H_{B,\alpha}u=(H_{B,0}u^{-},H_{B,0}u^{+})$. Since $\mathcal{M}_{\alpha}$ is real and self–adjoint, the operator $H_{B,\alpha}$ is self–adjoint in $L^{2}(\mathbb{R}^{2})$.
We note that the construction is invariant under gauge transformations. If $A^{\prime}=A+\nabla\chi$ with $\chi$ smooth and real–valued, then the unitary operator $U_{\chi}u=e^{i\chi}u$ satisfies $U_{\chi}H_{B,\alpha}U_{\chi}^{-1}=H_{B,\alpha}$, while the boundary maps transform covariantly (see, e.g., [15, Sec. 15] for gauge covariance in the Schrödinger setting). Indeed, for $U_{\chi}u=e^{i\chi}u$ one has
$\partial_{\nu}^{A+\nabla\chi}(U_{\chi}u)]=e^{i\chi}\,[\partial_{\nu}^{A}u],\qquad\gamma_{\mathrm{av}}(U_{\chi}u)=e^{i\chi}\gamma_{\mathrm{av}}u,$
so the boundary condition $\partial_{\nu}^{A}u]=\alpha\,\gamma_{\mathrm{av}}u$ is preserved. Thus the class of penetrable–wall Hamiltonians and their spectral properties do not depend on the particular choice of gauge.
More generally, the above construction is valid for a smooth magnetic potential $A$ and for variable wall strength $\alpha\in L^{\infty}(S_{a};\mathbb{R})$. In this setting the Green’s identity, the definition of the boundary maps, and the Krein’s resolvent formula remain valid [12, 6, 5, 17], and the extension $H_{B,\alpha}$ defined by the boundary condition $\partial_{\nu}^{A}u]=\mathcal{M}_{\alpha}\,\gamma_{\mathrm{av}}u$ is self–adjoint in $L^{2}(\mathbb{R}^{2})$. In the following sections, however, we shall restrict ourselves to the case of a constant magnetic field $B>0$ and a constant coupling $\alpha$, which allows for a fully explicit diagonalization of the Weyl function.
Remark 2.
The case $\Theta=0$ corresponds to the free magnetic Laplacian, whereas the limit $\Theta\to\infty$ yields a hard wall with Dirichlet condition $\Gamma_{1}u=0$ on $S_{a}$.
In the following sections we shall restrict ourselves to the case of a circular wall $S_{a}$ with constant coupling $\alpha\in\mathbb{R}$, which permits an explicit diagonalization of the Weyl function.
3 Landau levels broadened by a penetrable–wall
In this section we analyze how a penetrable–wall modifies the spectral picture of the Landau Hamiltonian. We first examine the behavior of negative eigenvalues in the regime of small magnetic fields, which provides the bridge to the non-magnetic case $H_{0,\alpha}$. We then prove that the essential spectrum is preserved for all $B>0$, while the infinite multiplicity of each Landau level is destroyed and replaced by sequences of discrete eigenvalues converging to the accumulation points. The subsequent mode decomposition and semiclassical interpretation clarify the detailed structure of the broadened spectrum (see also the general perturbative picture near Landau accumulation points in [18, 16, 7]).
3.1 Persistence of the essential spectrum
We begin by recalling the unperturbed picture. For a constant magnetic field $B>0$, the spectrum of the free magnetic Laplacian $H_{B,0}$ consists only of the Landau levels $\Lambda_{n}=B(2n+1)$, each of which is an eigenvalue of infinite multiplicity.
Once a penetrable–wall is introduced along $S_{a}$, the Hamiltonian becomes $H_{B,\alpha}$ with domain given in (2.7). Since the perturbation is supported on a compact curve, it is relatively compact with respect to $H_{B,0}$, and hence the essential spectrum remains unchanged by Weyl’s theorem.
Proposition 3.
For every $\alpha\in L^{\infty}(S_{a};\mathbb{R})$ the essential spectrum of $H_{B,\alpha}$ coincides with that of the free operator:
$\sigma_{\mathrm{ess}}(H_{B,\alpha})=\{\Lambda_{n}:\,n\in\mathbb{N}_{0}\}.$
Proof.
Fix a nonreal $z\in\mathbb{C}\setminus\{\Lambda_{n}:\ n\in\mathbb{N}_{0}\}$. By Krein’s resolvent formula (2.3) from Section 2 (see [6, 17] for the abstract form) one has
$(H_{B,\alpha}-z)^{-1}-(H_{B,0}-z)^{-1}$ (3.1)
$=$ $\gamma_{B}(z)\,(\mathcal{M}_{\alpha}-M_{B}(z))^{-1}\,\gamma_{B}(\bar{z})^{\ast}.$
Here $\gamma_{B}(z):H^{-1/2}(S_{a})\to L^{2}(\mathbb{R}^{2})$ is the Poisson operator, $M_{B}(z):H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$ is the Weyl function, and $\mathcal{M}_{\alpha}:H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$ is the multiplication operator by $\alpha\in L^{\infty}(S_{a};\mathbb{R})$, given explicitly in (2.6). For fixed nonreal $z$, the operator $(\mathcal{M}_{\alpha}-M_{B}(z))^{-1}$ is bounded by the general boundary triple theory.
We claim that $\gamma_{B}(\bar{z})^{\ast}:L^{2}(\mathbb{R}^{2})\to H^{1/2}(S_{a})$ is compact. Let $\chi\in C_{c}^{\infty}(\mathbb{R}^{2})$ be a cutoff that equals $1$ in a neighborhood of $S_{a}$. By elliptic regularity for magnetic Schrödinger operators with smooth $A$, the resolvent $(H_{B,0}-\bar{z})^{-1}$ maps $L^{2}(\mathbb{R}^{2})$ boundedly into $H^{2}_{\mathrm{loc}}(\mathbb{R}^{2})$. Hence, for any bounded domain $\Omega$ containing $S_{a}$, one has
$(H_{B,0}-\bar{z})^{-1}:L^{2}(\mathbb{R}^{2})\longrightarrow H^{2}(\Omega).$
The trace theorem for smooth boundaries then yields a bounded map
$H^{2}(\Omega)\ \xrightarrow{\ \mathrm{Tr}\ }\ H^{3/2}(S_{a}).$
Since $S_{a}$ is a compact 11–dimensional manifold, the Rellich–Kondrachov theorem ensures that
$H^{3/2}(S_{a})\hookrightarrow H^{1/2}(S_{a})$
is a compact embedding. This is a standard Sobolev result; see, for instance, [21, Ch. 4] or classical references on Sobolev spaces. Therefore the composition
$L^{2}(\mathbb{R}^{2})\ \xrightarrow{(H_{B,0}-\bar{z})^{-1}}\ H^{2}(\Omega)\ \xrightarrow{\ \mathrm{Tr}\ }\ H^{3/2}(S_{a})\ \hookrightarrow\ H^{1/2}(S_{a})$
is compact, showing that $\gamma_{B}(\bar{z})^{\ast}$ is compact.
Consequently, the right–hand side of (3.1) is a composition of bounded operators with the compact operator $\gamma_{B}(\bar{z})^{\ast}$, hence it is compact on $L^{2}(\mathbb{R}^{2})$. By Weyl’s theorem, compact perturbations do not change the essential spectrum. Since
$\sigma_{\mathrm{ess}}(H_{B,0})=\{\Lambda_{n}:\ n\in\mathbb{N}_{0}\},$
we conclude that
$\sigma_{\mathrm{ess}}(H_{B,\alpha})=\sigma_{\mathrm{ess}}(H_{B,0})=\{\Lambda_{n}:\ n\in\mathbb{N}_{0}\}.$
In particular, the Landau levels remain fixed as essential spectral points. $\Box$
3.2 Behavior of negative eigenvalues for small magnetic fields
We now discuss how the negative discrete spectrum of $H_{B,\alpha}$ behaves when a weak magnetic field $B>0$ is present. For $\alpha<0$ the operator $H_{0,\alpha}$ has finitely many negative eigenvalues below 0 (cf. [13, 1, 2]), which are separated from the essential spectrum $0,\infty)$. Our aim here is to show that these eigenvalues persist for small $B$, vary analytically with respect to $B$, and approach their non–magnetic counterparts as $B\to 0$.
The negative eigenvalues of $H_{0,\alpha}$ remain well below the lowest accumulation point $\Lambda_{0}=B$ for small $B$, and hence persist as isolated eigenvalues of $H_{B,\alpha}$. More precisely, Kato’s perturbation theory for analytic families of type (B) (see Kato [14, Chap. VII, Section 3.5]) ensures that every simple eigenvalue $E_{j}(0)<0$ of $H_{0,\alpha}$ generates an analytic eigenvalue branch $E_{j}(B)$ of $H_{B,\alpha}$ with $E_{j}(B)\to E_{j}(0)$ as $B\to 0$. Indeed, with form domain $\mathrm{dom}\,q_{B,\alpha}=H^{1}(\mathbb{R}^{2})$ the closed, lower–bounded quadratic form
$q_{B,\alpha}[u]:=\|(\nabla-iA_{B})u\|_{L^{2}(\mathbb{R}^{2})}^{2}+\alpha\,\|\gamma u\|_{L^{2}(S_{a})}^{2},\qquad A_{B}(x)=\tfrac{B}{2}(-x_{2},x_{1}),$
depends analytically on $B$. Hence $\{H_{B,\alpha}\}_{B\geq 0}$ is an analytic family of type (B) in the sense of Kato [14, Chap. VII, Section 3.5].
Moreover, by time-reversal symmetry the spectrum is invariant under $B\mapsto-B$, so the eigenvalue branches $E_{j}(B)$ are even functions of $B$ and the linear term in the expansion vanishes. The diamagnetic inequality [19, Eq. (1.1), Chap. 1] implies $E_{j}(B)\geq E_{j}(0)$, hence the quadratic coefficient is nonnegative. Therefore
$E_{j}(B)=E_{j}(0)+c_{j}B^{2}+o(B^{2}),\qquad c_{j}\geq 0.$
Thus the magnetic field has a stabilizing (non–lowering) effect on the negative eigenvalues, in the sense that they may shift upwards but cannot cross below their $B=0$ positions. Since the lowest Landau level $\Lambda_{0}=B$ tends to zero linearly in $B$, the gap between the negative eigenvalues and the essential spectrum remains open for sufficiently small $B$.
In summary, the negative spectrum of $H_{B,\alpha}$ for small $B$ is a perturbation of that of $H_{0,\alpha}$: the discrete eigenvalues survive, vary analytically with $B$, and approach their nonmagnetic counterparts as $B\to 0$. This behavior contrasts sharply with the spectral broadening near the Landau levels described in Sections 35, which have no analogue at $B=0$.
3.3 Spectral broadening and direction of accumulation
Although the Landau levels remain as essential spectral accumulation points, the penetrable–wall produces new families of discrete eigenvalues. These eigenvalues are not isolated once and for all, but rather they appear as sequences that converge to the Landau levels. Such broadening phenomena near Landau accumulation points are in line with the general perturbative picture (cf. [18, 16, 7]) and, for $\delta$–type interactions supported on curves, have been analyzed rigorously in [3].
Lemma 4 (Existence and uniqueness of mode solutions).
Fix $m\in\mathbb{Z}$ and $n\in\mathbb{N}_{0}$. Then the scalar equation
$\alpha-\mu_{m,B}(E)=0,\qquad E\in(\Lambda_{n},\Lambda_{n+1}),$ (3.2)
has at most one solution. Moreover, if $\alpha$ lies strictly between the boundary values $\lim_{E\downarrow\Lambda_{n}}\mu_{m,B}(E)$ and $\lim_{E\uparrow\Lambda_{n+1}}\mu_{m,B}(E)$, then such a solution exists.
Proof.
By Proposition 12, $\mu_{m,B}(E)$ is strictly monotone on each gap $(\Lambda_{n},\Lambda_{n+1})$. As $E\downarrow\Lambda_{n}$ one has $\mu_{m,B}(E)\to$-$\infty$, while as $E\uparrow\Lambda_{n+1}$ one has $\mu_{m,B}(E)\to+\infty$. Hence $\mu_{m,B}(E)$ covers the entire real line, and by the intermediate value theorem the scalar equation admits a unique solution whenever $\alpha$ lies between the endpoint values. $\Box$
Remark 3.
At special radii $a$ where $c_{n+1,m}=0$ (equivalently, when $L^{(|m|)}_{n+1}(\tfrac{B}{2}a^{2})=0$), the pole at $\Lambda_{n+1}$ disappears and $\mu_{m,B}(E)$ remains finite as $E\uparrow\Lambda_{n+1}$. In this exceptional case $\mu_{m,B}(E)$ ranges only over a half–line, so the existence of a solution depends on the sign of $\alpha$.
Theorem 5.
Assume that the wall is the circle $S_{a}$ and $\alpha\in\mathbb{R}\setminus\{0\}$. Then for every $n\in\mathbb{N}_{0}$ there exists an infinite sequence of eigenvalues of $H_{B,\alpha}$, obtained as solutions of the eigenvalue equations of Lemma 4, accumulating at $\Lambda_{n}$, except in the case $n=0$ with $\alpha>0$, where no such accumulation occurs. The accumulation is one–sided: if $\alpha<0$ the eigenvalues approach $\Lambda_{n}$ from above, while if $\alpha>0$ they approach from below (for $n=0$ this lower gap is absent).
Proof.
By Lemma 4, for each $m\in\mathbb{Z}$ the eigenvalue equation corresponding to the boundary condition in (2.7) has either no solution or a unique solution in $(\Lambda_{n},\Lambda_{n+1})$. From the expansion
$\mu_{m,B}(E)=\frac{c_{n,m}}{\Lambda_{n}-E}+h_{n,m}(E),$
with $c_{n,m}$ defined in (5.1), $c_{n,m}>0$, and $c_{n,m}\to 0$ as $|m|\to\infty$, it follows that for large $|m|$ a solution exists with
$E_{n,m}-\Lambda_{n}=-\frac{c_{n,m}}{\alpha}+o(c_{n,m}).$
Thus infinitely many $m$ contribute solutions converging to $\Lambda_{n}$, and the sign of the shift shows that the accumulation is from above for $\alpha<0$ and from below for $\alpha>0$. In particular, for $n\geq 1$ the eigenvalues originate from the upper gap $(\Lambda_{n},\Lambda_{n+1})$ when $\alpha<0$, and from the lower gap $(\Lambda_{n-1},\Lambda_{n})$ when $\alpha>0$. For $n=0$, only the case $\alpha<0$ yields eigenvalues converging to $\Lambda_{0}$ from above. $\Box$
Remark 4.
Although the infinite degeneracy of $\Lambda_{n}$ as an eigenvalue of $H_{B,0}$ is destroyed when $\alpha\neq 0$, the point $\Lambda_{n}$ itself remains in the essential spectrum of $H_{B,\alpha}$. This is not contradictory: the Landau levels cease to be eigenvalues of infinite multiplicity and instead become accumulation points of infinitely many discrete eigenvalues of finite multiplicity. Moreover, in the attractive case $\alpha<0$ the eigenvalues approach each Landau level from above, whereas for $\alpha>0$ no negative eigenvalues occur below $\Lambda_{0}=B$, and the broadened spectrum in each gap approaches $\Lambda_{n}$ from below.
Thus, the infinitely degenerate eigenvalue $\Lambda_{n}$ is replaced by a broadened spectrum lying in the neighboring gap. In other words, the infinitely degenerate point at $\Lambda_{n}$ is replaced by a countable family of eigenvalues accumulating at $\Lambda_{n}$, so that a whole sequence of discrete energies condenses towards the Landau level.
3.4 Mode decomposition in the circular case
When the wall is circular, $S_{a}$, and the coupling $\alpha$ is constant, the system enjoys rotational symmetry. This allows for a decomposition into angular momentum channels. Writing the Fourier basis $\{e^{im\theta}\}_{m\in\mathbb{Z}}$ on $S_{a}$, the Weyl function diagonalizes as
$M_{B}(z)e^{im\theta}=\mu_{m,B}(z)\,e^{im\theta},$
in accordance with the general boundary triple scheme and the special–function representation of Landau eigenfunctions (see [12, 6, 5] and, for explicit formulas, [8]).
Proposition 6.
For each $m\in\mathbb{Z}$ the eigenvalue condition reduces to the scalar equation
$\alpha-\mu_{m,B}(E)=0,\qquad E\in\mathbb{R}\setminus\{\Lambda_{n}\}.$
This equation admits at most one solution in each spectral gap. Collectively over all $m$, these solutions constitute the spectral broadening attached to $\Lambda_{n}$.
Proof.
Within the boundary triple framework fixed in Section 2 (see [12, 6, 5] for the abstract reduction and [17] for Krein-type formulas), we obtain the following characterization.
For $z\in\mathbb{C}\setminus\{\Lambda_{n}:\ n\in\mathbb{N}_{0}\}$ one has $\ker(T^{\ast}-z)\neq\{0\}$ and the Weyl function $M_{B}(z)$ is well defined on $H^{1/2}(S_{a})$. In particular, for real energies $E\in\mathbb{R}\setminus\{\Lambda_{n}\}$ one may evaluate $M_{B}(E)$ by restriction.
By the abstract eigenvalue characterization for boundary triples, $E$ is an eigenvalue of $H_{B,\alpha}$ if and only if there exists $\phi\in H^{1/2}(S_{a})\setminus\{0\}$ such that
$(\mathcal{M}_{\alpha}-M_{B}(E))\,\phi=0,$
where $\mathcal{M}_{\alpha}$ is the multiplication by the constant $\alpha$ on $H^{1/2}(S_{a})$ (cf. [12, 6]). In the circular case $S_{a}$, the operator $M_{B}(z)$ is diagonal in the Fourier basis $\{e^{im\theta}\}_{m\in\mathbb{Z}}$ of $L^{2}(S_{a})$, that is
$M_{B}(z)\,e^{im\theta}=\mu_{m,B}(z)\,e^{im\theta},\qquad m\in\mathbb{Z},$
with scalar coefficients $\mu_{m,B}(z)$. In particular, for real energies $E\in\mathbb{R}\setminus\{\Lambda_{n}\}$ one obtains
$M_{B}(E)\,e^{im\theta}=\mu_{m,B}(E)\,e^{im\theta}.$
Since $\mathcal{M}_{\alpha}$ commutes with $M_{B}(E)$ and, by (2.6), acts as $\mathcal{M}_{\alpha}e^{im\theta}=\alpha\,e^{im\theta}$, the eigenvalue condition reduces mode–wise to
$(\alpha-\mu_{m,B}(E))\,c_{m}=0\qquad(m\in\mathbb{Z}),$
where $c_{m}$ are the Fourier coefficients of $\phi=\sum_{m\in\mathbb{Z}}c_{m}e^{im\theta}$.
Hence a nontrivial solution $\phi$ exists if and only if there is an index $m$ such that
$\alpha-\mu_{m,B}(E)=0,$
which proves the stated scalar reduction.
It remains to show that there is at most one solution per gap. Fix $m\in\mathbb{Z}$. By Proposition 12 the map $E\mapsto\mu_{m,B}(E)$ is strictly monotone on every open interval $(\Lambda_{n},\Lambda_{n+1})$ (see also the general spectral monotonicity mechanisms for Landau-type operators in [18, 16]). Therefore the equation $\alpha-\mu_{m,B}(E)=0$ has at most one root in each gap. Since the full eigenfunction space is the orthogonal sum of the angular momentum channels, the discrete spectrum created by the wall is obtained by collecting, over all $m\in\mathbb{Z}$, the (gapwise unique, when present) solutions of these scalar equations. These eigenvalues, taken together as $m$ varies, form the cluster converging to $\Lambda_{n}$ described in Section 3. $\Box$
From this perspective, the degeneracy of the Landau level is broken into angular momentum sectors, and each sector contributes its own discrete eigenvalue. The location of the eigenvalue depends sensitively on the mode number $m$ and on the radius $a$ of the wall. Hence the distribution of eigenvalues in the broadened spectrum reflects the interplay between cyclotron motion and the geometry of the boundary.
To illustrate Proposition 6, we present in Figure 1 a numerical computation of the eigenvalues in the first spectral gap $(\Lambda_{0},\Lambda_{1})$ for $B=1.0$ and $a=1.1$. The discrete eigenvalues are shown as functions of the angular momentum index $m$, with different symbols corresponding to coupling strengths $\alpha=-0.5,-1.0,-1.5$. As predicted by the mode decomposition, each value of $m$ contributes at most one eigenvalue in the gap. Moreover, stronger attractive interaction (larger $|\alpha|$) pushes the eigenvalues further down into the gap, while for weaker coupling they cluster near the accumulation point $\Lambda_{1}=3.0$.
Figure 1: Eigenvalues in $(\Lambda_{0},\Lambda_{1})=(1.0,3.0)$ for $B=1.0$, $a=1.1$ (symbols: $\alpha=-0.5,-1.0,-1.5$). Dashed lines mark $\Lambda_{0},\Lambda_{1}$.
Remark 5 (Numerical remarks).
The eigenvalues displayed in Figures 13 were obtained by solving the scalar equation $\alpha-\mu_{m,B}(E)=0$ with a standard bracketing root finder. The strict monotonicity of $\mu_{m,B}$ on each spectral gap (Proposition 12) guarantees uniqueness and stability of the root when it exists. The series representation of $\mu_{m,B}$ was truncated according to the factorial/Laguerre estimates in Section 5, and convergence was checked by increasing the cutoffs. Only values stable under these checks are shown in the figures.
To further illustrate the effect of the penetrable–wall, Figure 2 compares the spatial profiles of the eigenfunctions in the free case ($\alpha=0$) and in the wall case ($\alpha=-3$), for $B=1$, $a=\sqrt{3}$, and angular momentum $m=1$. Both plots are $L^{2}$–normalized and share the same color scale. In the free case the Landau state is smoothly extended with a radial node near $r=2$, while in the presence of the attractive wall the wave function develops a clear localization around the boundary $r=a$. This visualization is consistent with the analytic conclusion that the wall interaction destroys the infinite degeneracy and creates localized eigenvalues accumulating at the Landau accumulation points.
Figure 2: Free ($\alpha=0$) vs. wall ($\alpha=-3$), with $B=1$, $a=\sqrt{3}$, $m=1$; both $L^{2}$-normalized, same color scale. The dashed circle marks the wall at $r=a$.
3.5 Special radii and embedded eigenvalues
In general the wall destroys the infinite multiplicity of each Landau level. Nevertheless, at certain discrete values of the radius $a$, remnants of the degeneracy can survive.
Lemma 7.
Fix $n\geq 1$.
  • (i)
    For each $m\in\mathbb{Z}$, the equation $L_{n}^{(|m|)}(x)=0$ has only finitely many positive solutions $x>0$ (hence the zero set in $x$ is discrete). Consequently, the set of radii $a>0$ such that $L_{n}^{(|m|)}\!\bigl{(}\tfrac{B}{2}a^{2}\bigr{)}=0$ is finite (hence discrete).
  • (ii)
    If $k\geq 1$ and the condition in (i) holds for both $m=k$ and m=km=-k, then the corresponding Landau eigenfunctions $\psi_{n,k}$ and $\psi_{n,-k}$ are linearly independent. Thus each such zero contributes multiplicity at most two to the eigenspace (for $k=0$ the contribution is at most one).
Proof of Lemma 7.
(i) Fix $m\in\mathbb{Z}$. The generalized Laguerre polynomial $L_{n}^{(|m|)}(x)$ is a polynomial in $x$ of degree $n$. Hence it has at most $n$ real zeros, in particular only finitely many positive zeros $x>0$. Therefore the zero set in $x$ is discrete. Since $x=\tfrac{B}{2}a^{2}$ is a strictly increasing $C^{\infty}$–bijection from $a>0$ onto $x>0$, the set of radii $a>0$ with $L_{n}^{(|m|)}\!\bigl{(}\tfrac{B}{2}a^{2}\bigr{)}=0$ is also discrete (indeed, finite, with cardinality $\leq n$).
(ii) Let $k\geq 1$ and suppose there exists $x>0$ such that $L_{n}^{(k)}(x)=0$ (equivalently, the condition in (i) holds for both $m=k$ and m=km=-k). Then the corresponding Landau eigenfunctions $\psi_{n,k}$ and $\psi_{n,-k}$ are linearly independent. For $k=0$ there is only one mode, so the contribution is at most one. Thus each such zero contributes multiplicity at most two to the eigenspace. $\Box$
Lemma 8.
Fix $n\geq 1$ and $x>0$. Then the set $\{\,m\in\mathbb{N}_{0}:L_{n}^{(m)}(x)=0\,\}$ has cardinality at most $n$.
Proof.
Using the finite expansion (see, e.g., standard formulas for generalized Laguerre polynomials),
$L_{n}^{(m)}(x)=\sum_{j=0}^{n}\frac{(-1)^{j}}{j!}\binom{n+m}{\,n-j\,}x^{j},$
and fix $x>0$. For each $j$, the binomial coefficient $\binom{n+m}{n-j}$ is a polynomial in $m$ of degree njn-j with leading term $m^{\,n-j}/(n-j)!$. Hence $L_{n}^{(m)}(x)$ is a polynomial in $m$ of exact degree $n$; indeed, the coefficient of $m^{n}$ comes solely from $j=0$ and equals $1/n!>0$, so the polynomial is not identically zero. Therefore it can have at most $n$ real zeros, and in particular at most $n$ integers $m\geq 0$ satisfy $L_{n}^{(m)}(x)=0$. $\Box$
Proposition 9.
For $n=0$ and $\alpha\neq 0$ one has $\ker(H_{B,\alpha}-\Lambda_{0})=\{0\}$. For $n\geq 1$ there exists a discrete set of radii $a$ such that $\ker(H_{B,\alpha}-\Lambda_{n})$ is nontrivial, with dimension not exceeding $2n$.
Proof.
We adopt the symmetric gauge $A(x)=\tfrac{B}{2}(-x_{2},x_{1})$, for which the radial component vanishes on $S_{a}$ ($A_{r}\equiv 0$). Hence the magnetic normal derivative coincides with the ordinary radial derivative on $S_{a}$, i.e., $\partial_{\nu}^{A}=\partial_{r}$. We first consider the lowest Landau level $n=0$. In the symmetric gauge one has
$\psi_{0,m}(r,\theta)=C_{0,m}\,r^{|m|}\,e^{-\tfrac{B}{4}r^{2}}\,e^{im\theta},$
so that $L_{0}^{(|m|)}\!\bigl{(}\tfrac{B}{2}r^{2}\bigr{)}\equiv 1$. Hence
$\psi_{0,m}(a,\theta)=C_{0,m}\,a^{|m|}\,e^{-\tfrac{B}{4}a^{2}}\,e^{im\theta}\neq 0\qquad\text{for all }m\in\mathbb{Z}.$
If $u\in\ker(H_{B,\alpha}-\Lambda_{0})$, then away from $S_{a}$ it solves $(H_{B,0}-\Lambda_{0})u=0$, so $u$ is a linear combination of lowest Landau eigenfunctions on each side of $S_{a}$. The penetrable–wall boundary condition, as in (2.7), is $\partial_{\nu}^{A}u]=\alpha,\gamma u$ on $S_{a}$.
Since $\psi_{0,m}(a,\theta)\neq 0$ for all $m$, the right–hand side $\alpha\,\gamma u$ can vanish only if $\gamma u=0$, which forces $u|_{S_{a}}\equiv 0$. Unique continuation then implies $u\equiv 0$. Therefore $\ker(H_{B,\alpha}-\Lambda_{0})=\{0\}$.
Now let $n\geq 1$. If $L_{n}^{(|m|)}\!\bigl{(}\tfrac{B}{2}a^{2}\bigr{)}=0$, then $\psi_{n,m}$ vanishes identically on $S_{a}$ and thus belongs to $\ker(H_{B,\alpha}-\Lambda_{n})$. By Lemma 7(i), the set of such radii $a$ is discrete. Moreover, Lemma 8 shows that for fixed $n$ there are at most $n$ indices $m\geq 0$ for which the condition can be satisfied. Taking into account the symmetry $m\mapsto-m$ and Lemma 7(ii), each such index contributes at most two independent eigenfunctions $\psi_{n,\pm m}$, so the multiplicity of the eigenspace does not exceed $2n$. $\Box$
Thus, embedded eigenvalues at the Landau accumulation points may persist, but only in a highly nongeneric manner and with strictly limited multiplicity.
3.6 Semiclassical picture of mode resonance
For intuition, we compare this description with the classical cyclotron picture. For an electron in the nn–th Landau level, the cyclotron radius is given by
$r_{c}(n)\approx\sqrt{2n+1}\,\ell_{B},$ (3.3)
where $\ell_{B}=1/\sqrt{B}$ denotes the magnetic length (see, e.g., [15]). This approximation reflects the classical orbit size associated with the nn–th Landau state.
In the mode decomposition, the mm–th angular momentum channel of a Landau eigenfunction has its radial probability concentrated near $r\approx\sqrt{2m/B}$. Thus when the wall radius $a$ is close to $r_{c}(n)$ in (3.3), there exists a mode index
$m_{\ast}\approx\frac{a^{2}}{2\ell_{B}^{2}}-\left(n+\frac{1}{2}\right),$
whose cyclotron orbit is resonant with the wall position. For this $m_{\ast}$–mode, the boundary interaction acts most strongly, so that the associated eigenvalue $E_{n,m_{\ast}}$ deviates from $\Lambda_{n}$ more significantly than for neighboring modes. Consequently, the point $E_{n,m_{\ast}}$ appears as the most prominent element in the cluster of eigenvalues condensing at $\Lambda_{n}$.
As the wall radius $a$ varies, the resonant index $m_{\ast}$ changes, and the “leading” eigenvalue in the cluster jumps from one mode to another. This semiclassical resonance mechanism explains how the wall geometry governs the fine structure of the broadened spectrum. We stress that the word “resonance” here is used in a heuristic sense, referring to the geometric matching between the wall and the cyclotron orbit, not to resonances in the spectral or scattering theoretic sense. See also the general perturbative picture near Landau accumulation points in [18, 16, 7].
Figure 3 shows the radial probability distribution $2\pi r|\psi_{n,m}(r)|^{2}$ for $n=0$ and angular momentum indices $m=3,4$, normalized to unit height. The vertical dashed line indicates the wall at $a=3.0$. One observes that the $m=3$ mode reaches its maximum almost exactly at the wall, while the $m=4$ mode peaks slightly outside. These neighboring modes therefore have the largest overlap with the boundary, so the wall interaction affects them most strongly. Consequently their eigenvalue shifts are more pronounced than those of other modes, accounting for their prominence in the broadened spectrum.
Figure 3: Radial profiles ($n=0$, $m=3,4$) for $B=1$, $a=3$; normalized to unit height. Dashed line: $r=a$.
4 Explicit form of the Weyl function
In this section we derive an explicit representation of the Weyl function, which will serve as the basis for the monotonicity and asymptotic analysis in the subsequent section.
In order to make the analysis concrete it is convenient to write down an explicit expression for the diagonal entries of the Weyl function. We take the symmetric gauge $A(x)=\tfrac{B}{2}(-x_{2},x_{1})$, in which the Landau eigenfunctions are well known (see e.g., [15]). For integers $m\in\mathbb{Z}$ and $n\in\mathbb{N}_{0}$, an orthonormal basis of the eigenspace $\ker(H_{B,0}-\Lambda_{n})$ with $\Lambda_{n}=B(2n+1)$ is given by
$\psi_{n,m}(r,\theta)=C_{n,m}\,r^{|m|}\,e^{-\tfrac{B}{4}r^{2}}\,L_{n}^{(|m|)}\Bigl{(}\tfrac{B}{2}r^{2}\Bigr{)}\,e^{im\theta},$
where $(r,\theta)$ are polar coordinates, $L_{n}^{(|m|)}$ is the generalized Laguerre polynomial (standard properties can be found in [8]), and $C_{n,m}$ is a normalization constant. These eigenfunctions reflect the rotational symmetry and they separate the angular momentum channels.
When restricted to the circle of radius $a$, these functions take the form
$\psi_{n,m}(a,\theta)=C_{n,m}\,a^{|m|}\,e^{-\tfrac{B}{4}a^{2}}\,L_{n}^{(|m|)}\Bigl{(}\tfrac{B}{2}a^{2}\Bigr{)}\,e^{im\theta}.$
Proposition 10 (Identification of the Weyl function).
Let $M_{B}(z):H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$ be the Weyl function of the (quasi) boundary triple from Section 2.3, and define
$K_{B}(z):=\tau\,(H_{B,0}-z)^{-1}\,\tau^{\ast}\quad\text{on }L^{2}(S_{a}).$
Then, for all $\varphi,\psi\in H^{1/2}(S_{a})$,
$\langle M_{B}(z)\varphi,\psi\rangle_{H^{-1/2},H^{1/2}}=\langle K_{B}(z)\varphi,\psi\rangle_{L^{2}(S_{a})}.$
In particular, identifying $H^{1/2}(S_{a})\hookrightarrow L^{2}(S_{a})\hookrightarrow H^{-1/2}(S_{a})$ via the $L^{2}$–pivot, $M_{B}(z)$ and $K_{B}(z)$ agree as operators $H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$.
Proof.
Fix $z\in\mathbb{C}\setminus\{\Lambda_{n}:\,n\in\mathbb{N}_{0}\}$. By the (quasi) boundary triple construction (Section 2.3), for $\varphi\in H^{1/2}(S_{a})$ let $u=\gamma_{B}(z)\varphi\in\ker(T^{\ast}-z)$ with $\Gamma_{0}u=\varphi$ and $M_{B}(z)\varphi=\Gamma_{1}u$. For $\psi\in H^{1/2}(S_{a})$ set $v=\gamma_{B}(\bar{z})\psi\in\ker(T^{\ast}-\bar{z})$. Green’s identity (Section 2.2) yields
$(z-\bar{z})\langle u,v\rangle_{L^{2}(\mathbb{R}^{2})}=\langle\varphi,\Gamma_{1}v\rangle_{H^{-1/2},H^{1/2}}-\langle\Gamma_{1}u,\psi\rangle_{H^{1/2},H^{-1/2}}.$
Since $\Gamma_{1}$ is the average trace on $S_{a}$ and $\tau$ is the trace to $S_{a}$, we have $\langle\varphi,\Gamma_{1}v\rangle_{H^{-1/2},H^{1/2}}=\langle\varphi,\tau v\rangle_{L^{2}(S_{a})}$ and $\langle\Gamma_{1}u,\psi\rangle_{H^{1/2},H^{-1/2}}=\langle\tau u,\psi\rangle_{L^{2}(S_{a})}$. On the other hand, with $w:=(H_{B,0}-z)^{-1}\tau^{\ast}\varphi$ we have $(H_{B,0}-z)w=\tau^{\ast}\varphi$; by the Poisson construction and uniqueness, $\tau u=\tau w$, and similarly for $v$ with $\bar{z}$. Therefore
$\langle M_{B}(z)\varphi,\psi\rangle_{H^{-1/2},H^{1/2}}$ $=$ $\langle\tau u,\psi\rangle_{L^{2}(S_{a})}$
$=$ $\langle\tau(H_{B,0}-z)^{-1}\tau^{\ast}\varphi,\psi\rangle_{L^{2}(S_{a})}$
$=$ $\langle K_{B}(z)\varphi,\psi\rangle_{L^{2}(S_{a})},$
where $K_{B}(z):=\tau(H_{B,0}-z)^{-1}\tau^{\ast}$ on $L^{2}(S_{a})$. This proves the claim. $\Box$
Lemma 11 (Diagonalization of $K_{B}(z)$).
For each $m\in\mathbb{Z}$, the operator
$K_{B}(z):=\tau\,(H_{B,0}-z)^{-1}\,\tau^{\ast}\quad\text{on }L^{2}(S_{a})$
is diagonal in the Fourier basis $\{e^{im\theta}\}_{m\in\mathbb{Z}}$ of $L^{2}(S_{a})$, and
$K_{B}(z)\,e^{im\theta}=\mu_{m,B}(z)\,e^{im\theta},$
with coefficients
$\mu_{m,B}(z)=\sum_{n=0}^{\infty}\frac{\|\psi_{n,m}(a,\cdot)\|_{L^{2}(S_{a})}^{2}}{\Lambda_{n}-z},\qquad z\in\mathbb{C}\setminus\{\Lambda_{n}:\,n\in\mathbb{N}_{0}\}.$ (4.1)
Proof.
Let $R_{B}(z)=(H_{B,0}-z)^{-1}$ for $z\in\mathbb{C}\setminus\{\Lambda_{n}:\,n\in\mathbb{N}_{0}\}$ and let $\tau$ be the trace map to $S_{a}$, so that $\tau u=u|_{S_{a}}$ for smooth $u$. We consider the boundary operator
$K_{B}(z):=\tau\,R_{B}(z)\,\tau^{\ast}\quad\text{on }L^{2}(S_{a}).$
By Proposition 10 we identify $K_{B}(z)$ with the Weyl function $M_{B}(z):H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$ in the sense that
$\langle M_{B}(z)\varphi,\psi\rangle_{H^{-1/2},H^{1/2}}=\langle K_{B}(z)\varphi,\psi\rangle_{L^{2}(S_{a})}\quad(\varphi,\psi\in H^{1/2}(S_{a})).$
Accordingly, it suffices to diagonalize $K_{B}(z)$ on $L^{2}(S_{a})$ and then read off the diagonal of $M_{B}(z)$. In the present setting $K_{B}(z)$ agrees with the Weyl function $M_{B}(z)$, since the Poisson solution operator and the jump map are related to $R_{B}(z)$ by Green’s identity (2.2) and the choice of boundary maps. Thus it is enough to compute $K_{B}(z)$.
The free Hamiltonian $H_{B,0}$ is invariant under planar rotations about the origin, and the same is true for $\tau$ and $\tau^{\ast}$. Hence $K_{B}(z)$ commutes with the unitary representation of the rotation group on $L^{2}(S_{a})$. The irreducible subspaces of this representation are the one–dimensional spaces spanned by $e^{im\theta}$, $m\in\mathbb{Z}$. It follows that $K_{B}(z)$ is diagonal in the basis $\{e^{im\theta}\}_{m\in\mathbb{Z}}$ and acts as a scalar on each mode. Therefore there exist numbers $\mu_{m,B}(z)$ such that
$K_{B}(z)\,e^{im\theta}=\mu_{m,B}(z)\,e^{im\theta},\qquad m\in\mathbb{Z}.$
To identify $\mu_{m,B}(z)$ we use the spectral resolution of $R_{B}(z)$. Let $\{\psi_{n,k}\}_{n\in\mathbb{N}_{0},\ k\in\mathbb{Z}}$ be an orthonormal basis of Landau eigenfunctions in the symmetric gauge (cf. [15, Sec. 15]), so that
$R_{B}(z)=\sum_{n=0}^{\infty}\frac{\Pi_{n}}{\Lambda_{n}-z},\qquad\Pi_{n}=\sum_{k\in\mathbb{Z}}|\psi_{n,k}\rangle\langle\psi_{n,k}|,\qquad\Lambda_{n}=B(2n+1),$
with convergence in the strong operator topology on $L^{2}(\mathbb{R}^{2})$. Then
$K_{B}(z)\,e^{im\theta}$ $=$ $\tau\,R_{B}(z)\,\tau^{\ast}e^{im\theta}=\sum_{n=0}^{\infty}\frac{1}{\Lambda_{n}-z}\,\tau\,\Pi_{n}\,\tau^{\ast}e^{im\theta}$ (4.2)
$=$ $\sum_{n=0}^{\infty}\frac{1}{\Lambda_{n}-z}\sum_{k\in\mathbb{Z}}\tau\psi_{n,k}\,\langle\psi_{n,k},\tau^{\ast}e^{im\theta}\rangle.$
By definition of the adjoint, $\langle\psi_{n,k},\tau^{\ast}f\rangle_{L^{2}(\mathbb{R}^{2})}=\langle\tau\psi_{n,k},f\rangle_{L^{2}(S_{a})}$ for $f\in L^{2}(S_{a})$. On the circle $S_{a}$ one has
$\tau\psi_{n,k}(a,\theta)=C_{n,k}\,a^{|k|}e^{-\tfrac{B}{4}a^{2}}L_{n}^{(|k|)}\!\left(\tfrac{B}{2}a^{2}\right)e^{ik\theta}=:A_{n,k}e^{ik\theta}.$
Consequently,
$\|\psi_{n,m}(a,\cdot)\|^{2}_{L^{2}(S_{a})}=\int_{S_{a}}|\tau\psi_{n,m}|^{2}\,d\sigma=\int_{0}^{2\pi}|A_{n,m}e^{im\theta}|^{2}\,a\,d\theta=2\pi a|A_{n,m}|^{2},$ (4.3)
which is consistent with the definition of $c_{n,m}$ in (5.1). Hence the orthogonality of $\{e^{ik\theta}\}$ yields
$\langle\tau\psi_{n,k},e^{im\theta}\rangle_{L^{2}(S_{a})}=\begin{cases}2\pi aA_{n,m},&k=m,\\[6.0pt] 0,&k\neq m.\end{cases}$
Hence only the terms with $k=m$ survive in (4.2), and we find
$K_{B}(z)\,e^{im\theta}$ $=$ $\sum_{n=0}^{\infty}\frac{1}{\Lambda_{n}-z}\,\Big{(}\tau\psi_{n,m}\Big{)}\,\langle\tau\psi_{n,m},e^{im\theta}\rangle_{L^{2}(S_{a})}$ (4.4)
$=$ $\sum_{n=0}^{\infty}\frac{1}{\Lambda_{n}-z}\,\Big{(}A_{n,m}e^{im\theta}\Big{)}\,\Big{(}2\pi a\overline{A_{n,m}}\Big{)}$
$=$ $\Bigg{(}\sum_{n=0}^{\infty}\frac{2\pi a|A_{n,m}|^{2}}{\Lambda_{n}-z}\Bigg{)}\,e^{im\theta}.$
Since
$\|\psi_{n,m}(a,\cdot)\|_{L^{2}(S_{a})}^{2}=\int_{S_{a}}|\tau\psi_{n,m}|^{2}\,d\sigma=\int_{0}^{2\pi}|A_{n,m}|^{2}\,a\,d\theta=2\pi a|A_{n,m}|^{2},$
the coefficient in (4.4) equals $\|\psi_{n,m}(a,\cdot)\|_{L^{2}(S_{a})}^{2}$ by (4.3), so that
$\mu_{m,B}(z)=\sum_{n=0}^{\infty}\frac{\|\psi_{n,m}(a,\cdot)\|_{L^{2}(S_{a})}^{2}}{\Lambda_{n}-z}.$
Finally, inserting the explicit expression for $A_{n,m}$ gives
$\|\psi_{n,m}(a,\cdot)\|_{L^{2}(S_{a})}^{2}=2\pi a\,|C_{n,m}|^{2}\,a^{2|m|}\,e^{-\tfrac{B}{2}a^{2}}\left|L_{n}^{(|m|)}\Big{(}\tfrac{B}{2}a^{2}\Big{)}\right|^{2},$
and therefore
$\mu_{m,B}(z)=2\pi a\,e^{-\tfrac{B}{2}a^{2}}\sum_{n=0}^{\infty}\frac{|C_{n,m}|^{2}\,a^{2|m|}\,\big{|}L_{n}^{(|m|)}(\tfrac{B}{2}a^{2})\big{|}^{2}}{\Lambda_{n}-z}.$
This proves the stated formulas. $\Box$
Remark 6.
The normalization constants can be chosen as
$|C_{n,m}|^{2}=\frac{B^{|m|+1}\,n!}{2^{|m|+1}\pi\,(n+|m|)!},$
and hence the coefficients $\mu_{m,B}(z)$ admit a completely explicit representation.
Proposition 12 (Monotonicity of the diagonal coefficients).
For each fixed $m\in\mathbb{Z}$, the function $E\mapsto\mu_{m,B}(E)$ is strictly monotone on every open interval $(\Lambda_{n},\Lambda_{n+1})$. It follows that the eigenvalue equation
$\alpha-\mu_{m,B}(E)=0$
has at most one solution in each spectral gap for every angular momentum mode. Thus each mode contributes at most one discrete eigenvalue, and the entire spectral broadening attached to a Landau level is obtained by collecting these contributions over all integers $m$.
Proof.
Let $R_{B}(z)=(H_{B,0}-z)^{-1}$ for $z\in\mathbb{C}\setminus\{\Lambda_{n}:\,n\in\mathbb{N}_{0}\}$ and let $\tau$ denote the trace to $S_{a}$. By Proposition 10, the Weyl function $M_{B}(z)$ may be identified with $K_{B}(z)=\tau R_{B}(z)\tau^{\ast}$ on $L^{2}(S_{a})$. As noted in (2.4), one has
$M_{B}(z)=\tau R_{B}(z)\tau^{\ast},\qquad\mu_{m,B}(z)=\langle e^{im\theta},\,M_{B}(z)\,e^{im\theta}\rangle_{L^{2}(S_{a})}.$
Fix $m\in\mathbb{Z}$ and let $E\in(\Lambda_{n},\Lambda_{n+1})$. Since $E$ belongs to the resolvent set of $H_{B,0}$, the map $z\mapsto M_{B}(z)$ is analytic near $E$, and the resolvent identity gives
$\frac{d}{dE}R_{B}(E)=(H_{B,0}-E)^{-2}.$
Differentiating $\mu_{m,B}(E)=\langle e^{im\theta},\,\tau R_{B}(E)\tau^{\ast}e^{im\theta}\rangle$ yields
$\frac{d}{dE}\mu_{m,B}(E)$ $=\Big{\langle}e^{im\theta},\,\tau(H_{B,0}-E)^{-2}\tau^{\ast}e^{im\theta}\Big{\rangle}_{L^{2}(S_{a})}$
$=\Big{\langle}(H_{B,0}-E)^{-1}\tau^{\ast}e^{im\theta},\,(H_{B,0}-E)^{-1}\tau^{\ast}e^{im\theta}\Big{\rangle}_{L^{2}(\mathbb{R}^{2})}$
$=\Big{\|}(H_{B,0}-E)^{-1}\tau^{\ast}e^{im\theta}\Big{\|}_{L^{2}(\mathbb{R}^{2})}^{2}.$ (4.5)
Here we used that (cf. (2.4))
$\tau(H_{B,0}-E)^{-2}\tau^{\ast}=\bigl{(}(H_{B,0}-E)^{-1}\tau^{\ast}\bigr{)}^{\ast}\bigl{(}(H_{B,0}-E)^{-1}\tau^{\ast}\bigr{)},$
which is a positive operator on $L^{2}(S_{a})$. Therefore the right–hand side of (4.5) is nonnegative for all $E$ in the open gaps $(\Lambda_{n},\Lambda_{n+1})$. In fact it is strictly positive: indeed, $\tau^{\ast}e^{im\theta}\neq 0$ in $L^{2}(\mathbb{R}^{2})$ (since $e^{im\theta}$ is a nontrivial boundary function), and the free resolvent $(H_{B,0}-E)^{-1}$ is injective on $L^{2}(\mathbb{R}^{2})$ whenever $E\notin\{\Lambda_{n}:\,n\in\mathbb{N}_{0}\}$. Hence the norm in (4.5) cannot vanish, and $E\mapsto\mu_{m,B}(E)$ is strictly increasing on every open spectral gap. It follows that for each fixed $\alpha\in\mathbb{R}$ the function $E\mapsto\alpha-\mu_{m,B}(E)$ is strictly monotone on every gap, so the equation $\alpha-\mu_{m,B}(E)=0$ admits at most one solution there. $\Box$
5 Asymptotic behavior of boundary coefficients
In this section we investigate the asymptotic behavior of the boundary coefficients for large $|m|$, which provides the basis for the precise evaluation of the eigenvalue shifts.
We study the large $|m|$ behavior of the boundary coefficients
$c_{n,m}=\|\psi_{n,m}(a,\cdot)\|_{L^{2}(S_{a})}^{2}=2\pi a\,|C_{n,m}|^{2}\,a^{2|m|}\,e^{-\tfrac{B}{2}a^{2}}\Big{|}L_{n}^{(|m|)}\Big{(}\tfrac{B}{2}a^{2}\Big{)}\Big{|}^{2},$ (5.1)
where $|C_{n,m}|^{2}=\dfrac{B^{|m|+1}\,n!}{2^{|m|+1}\pi\,(n+|m|)!}$. The estimates below rely on standard asymptotics for factorials and generalized Laguerre polynomials; see, e.g., [8].
Lemma 13.
Fix $n\in\mathbb{N}_{0}$, $a>0$, $B>0$. As $|m|\to\infty$ one has
$|C_{n,m}|^{2}\sim\frac{1}{\pi}\,\frac{B^{|m|+1}\,n!}{2^{|m|+1}}\,\frac{e^{\,n+|m|}}{(n+|m|)^{\,n+|m|+1/2}}.$
In particular, for some constant $K_{1}=K_{1}(n,a,B)>0$,
$|C_{n,m}|^{2}\leq K_{1}\,\frac{1}{(n+|m|)!}\,\Big{(}\frac{eB}{2}\Big{)}^{|m|}\qquad\text{for all sufficiently large }|m|.$
Proof.
Stirling’s formula gives
$(n+|m|)!\sim\sqrt{2\pi(n+|m|)}\,\Big{(}\frac{n+|m|}{e}\Big{)}^{n+|m|},\qquad|m|\to\infty,$
and inserting this in the definition of $|C_{n,m}|^{2}$ yields the claimed asymptotics. The one–sided bound follows from the two–sided Stirling estimates by absorbing constants into $K_{1}$ (see, e.g., [8, Chap. 1] for a convenient reference). $\Box$
Lemma 14.
Fix $n\in\mathbb{N}_{0}$ and $x>0$. As $k\to\infty$,
$L_{n}^{(k)}(x)=\binom{n+k}{n}\Big{(}1-\frac{x}{k}+O(k^{-2})\Big{)}.$
In particular, there exists $K_{2}=K_{2}(n,x)>0$ such that
$\Big{|}L_{n}^{(k)}(x)\Big{|}\leq K_{2}\,\binom{n+k}{n}\qquad\text{for all large }k.$
Proof.
Using the finite expansion
$L_{n}^{(k)}(x)=\sum_{j=0}^{n}\frac{(-1)^{j}}{j!}\binom{n+k}{n-j}\,x^{\,j},$
together with the elementary binomial estimate $\binom{n+k}{n-j}=\binom{n+k}{n}\frac{n!}{(n-j)!}\,k^{-j}\big{(}1+O(k^{-1})\big{)}$ for fixed $n$, we obtain the stated asymptotics; see, e.g., [8, Chap. 10]. The bound follows by taking absolute values for large $k$ and absorbing constants. $\Box$
Lemma 15.
Fix $n\in\mathbb{N}_{0}$. For $E\in\mathbb{R}$ define
$h_{n,m}(E):=\sum_{k\neq n}\frac{c_{k,m}}{\Lambda_{k}-E},$
where the summand with $k=n$ is omitted. In particular, $h_{n,m}(\Lambda_{n})$ denotes the same series evaluated at $E=\Lambda_{n}$, which is well defined since $\sum_{k\geq 0}c_{k,m}<\infty$ uniformly in $m$ and $|\Lambda_{k}-\Lambda_{n}|=2B|k-n|$. Then $h_{n,m}(\Lambda_{n})\to 0$ as $|m|\to\infty$. Moreover, the convergence is uniform: for every $\varepsilon>0$ there exists $M(\varepsilon)$ such that $|h_{n,m}(\Lambda_{n})|<\varepsilon$ whenever $|m|\geq M(\varepsilon)$.
Proof.
Since $|\Lambda_{k}-\Lambda_{n}|=2B\,|k-n|$, we have
$|h_{n,m}(\Lambda_{n})|\leq\frac{1}{2B}\sum_{k\neq n}\frac{c_{k,m}}{|k-n|}.$
Split the sum at an arbitrary $J\in\mathbb{N}$:
$|h_{n,m}(\Lambda_{n})|\leq\frac{1}{2B}\sum_{|k-n|\leq J}\frac{c_{k,m}}{|k-n|}\;+\;\frac{1}{2BJ}\sum_{k\geq 0}c_{k,m}.$
For the finite head, each fixed $k$ satisfies $c_{k,m}\to 0$ as $|m|\to\infty$ (by Proposition 16, $c_{k,m}$ decays faster than any exponential in $|m|$), hence the head tends to 0 as $|m|\to\infty$. For the tail, standard kernel and generating–function estimates for generalized Laguerre polynomials, together with the explicit normalization (cf. Remark 6), yield a constant $C=C(a,B)>0$ such that
$\sum_{k\geq 0}c_{k,m}\leq C\qquad\text{for all }m\in\mathbb{Z}.$
Indeed, one can use the generating function identity
$\sum_{k=0}^{\infty}L_{k}^{(|m|)}(r)^{2}\,t^{k}=\frac{1}{1-t}\exp\left(-\frac{tr}{1-t}\right)I_{|m|}\left(\frac{2\sqrt{tr}}{1-t}\right),\qquad|t|<1,$
(cf. [20]) to deduce such a bound uniformly in $m$. Thus the tail is bounded by $C/(2BJ)$. First choose $J$ large to make the tail $\varepsilon/2$, then choose $|m|$ large so that the head $\varepsilon/2$. This proves the claim. $\Box$
Proposition 16.
Fix $n\in\mathbb{N}_{0}$, $a>0$, $B>0$. Set $x=\tfrac{B}{2}a^{2}$. Then, as $|m|\to\infty$,
$c_{n,m}\sim K_{n}(a,B)\,\frac{1}{(n+|m|)!}\,\Big{(}\frac{Ba^{2}}{2}\Big{)}^{|m|}\,\binom{n+|m|}{n}^{\!2},$
for some positive constant $K_{n}(a,B)$. In particular, $c_{n,m}$ decays faster than any exponential in $|m|$. (That is, for every $\beta>0$ one has $c_{n,m}=O(e^{-\beta|m|})$ as $|m|\to\infty$.)
Proof.
Combining the defining formula (5.1) for $c_{n,m}$ with the expression for $\|\psi_{n,m}(a,\cdot)\|^{2}_{L^{2}(S_{a})}$ in (4.3) and the estimates of the previous two lemmas gives
$c_{n,m}$ $=$ $2\pi a\,|C_{n,m}|^{2}\,a^{2|m|}\,e^{-\tfrac{B}{2}a^{2}}\Big{|}L_{n}^{(|m|)}(x)\Big{|}^{2}$ (5.2)
$\sim$ $\Big{(}2\pi a\,e^{-\tfrac{B}{2}a^{2}}\Big{)}\,\frac{B^{|m|+1}\,n!}{2^{|m|+1}\pi\,(n+|m|)!}\,a^{2|m|}\,\binom{n+|m|}{n}^{\!2}$
$=$ $K_{n}(a,B)\,\frac{1}{(n+|m|)!}\,\Big{(}\frac{Ba^{2}}{2}\Big{)}^{|m|}\,\binom{n+|m|}{n}^{\!2},$
with $K_{n}(a,B)=n!Be^{-\tfrac{B}{2}a^{2}}/2$. Finally, using $\binom{n+|m|}{n}\sim(n+|m|)^{n}/n!$ for fixed $n$ and Stirling’s formula for $(n+|m|)!$, we obtain the stated faster than any exponential decay (see also [8, Chap. 10]). $\Box$
Theorem 17.
Let $n\in\mathbb{N}_{0}$ and $\alpha\in\mathbb{R}$. For the eigenvalues $E_{n,m}$ solving $\alpha-\mu_{m,B}(E)=0$ in a neighborhood of $\Lambda_{n}$, with $c_{n,m}$ defined in (5.1), one has
$E_{n,m}-\Lambda_{n}=-\frac{c_{n,m}}{\alpha}+o(c_{n,m})\qquad(|m|\to\infty).$
In particular, the shifts $E_{n,m}-\Lambda_{n}$ are negative for $\alpha>0$ and positive for $\alpha<0$, hence always approach $\Lambda_{n}$ from the side predicted in Theorem 5.
Proof.
Write
$\mu_{m,B}(E)=\frac{c_{n,m}}{\Lambda_{n}-E}+h_{n,m}(E),$
where $h_{n,m}$ is real–analytic near $\Lambda_{n}$, and $h_{n,m}(\Lambda_{n})=\sum_{k\neq n}\frac{c_{k,m}}{\Lambda_{k}-\Lambda_{n}}$. Set $E=\Lambda_{n}+\delta$. Then
$\alpha-\mu_{m,B}(\Lambda_{n}+\delta)=\alpha+\frac{c_{n,m}}{\delta}-h_{n,m}(\Lambda_{n})-r_{n,m}(\delta),\qquad r_{n,m}(\delta)\to 0\ (\,\delta\to 0\,).$
By Lemma 15, $h_{n,m}(\Lambda_{n})=o(1)$ uniformly as $|m|\to\infty$. Since $c_{n,m}\to 0$ faster than any exponential in $|m|$ by Proposition 16, for $|m|$ large we may assume $|h_{n,m}(\Lambda_{n})|\leq|\alpha|/4$ and $|r_{n,m}(\delta)|\leq|\alpha|/4$ at the (unique small) solution $\delta$. Solving for $\delta$ yields
$\delta=\frac{c_{n,m}}{h_{n,m}(\Lambda_{n})+r_{n,m}(\delta)-\alpha}=-\frac{c_{n,m}}{\alpha}\bigl{(}1+o(1)\bigr{)}\qquad(|m|\to\infty),$
and hence
$E_{n,m}-\Lambda_{n}=-\frac{c_{n,m}}{\alpha}+o(c_{n,m}).$
The sign statement follows immediately. $\Box$
Appendix: Numerical methods
The numerical plots in Figures 13 were obtained from the scalar eigenvalue condition (3.2) using the following standard procedures.
  • (i)
    Series truncation. The series representation of $\mu_{m,B}(z)$ in (4.1) was truncated at $n\leq N_{\max}$, chosen so that the last term was smaller than $10^{-12}$ in absolute value. Increasing $N_{\max}$ did not change the displayed digits of the results.
  • (ii)
    Root finding. For each angular momentum index $m$, the equation was solved in the relevant gap $(\Lambda_{n},\Lambda_{n+1})$ by the Brent method, starting from an interval bracketing the root. The strict monotonicity of $\mu_{m,B}(E)$ (Proposition 12) guarantees uniqueness in each gap, so a simple bracketing suffices.
  • (iii)
    Accuracy. Solutions are stable under refinement of the truncation and bracketing. Relative errors are estimated to be less than $10^{-8}$.
The figures were produced with standard numerical software (Python/SciPy). The computations are straightforward and can be reproduced with any comparable package.